Use the following financial information to find the entry you would make on an income statement for GROSS MARGIN for the year ended December 31, 2011: Gross Sales, $241,000; Sales Returns and Allowances, $7,500; Sales Discounts, $7,800; Merchandise Inventory, January 1, 2011, $45,200; Merchandise Inventory, December 31, 2011, $71,300; Net Purchases, $91,900; Freight In, $1950; Salaries, $97,000; Rent, $30,600; Utilities, $2,500; Insurance, $2,450; and Income Tax, $19,350.
Answer #1

Gross margin = Net sales - Cost of goods sold

Net sales = Gross Sales - Sales Returns and Allowances - Sales discounts

Net sales = 241,000 - 7,500 - 7,800 = 225,700

Cost of goods sold = Opening Stock + Net purchases + Freight in - Closing stock

Cost of goods sold = 45,200 + 91,900 + 1,950 - 71,300 = 67,750

Gross margin = 225,700 - 67,750 = $157,950